7 Graph the system of linear inequalities 5x+4y= 7) · 2x + 7y = -85
Answer #1

Given :- 5x+4y= 7 2x + 74 = -8 52 + 4y = 7 put x=0, then y=? 560) + 4y=7 y = 714 = 1.75 x=1, then y= 2 5(1) + 49 = 7 4y = 2 y(0, 1.75) (-4, 0) (1.4, 0) - -2 (0, -1.143) (3,-2)